The mean age of the 30 members of the Ft. Wayne Comets is 28. After the
current season, five members whose mean age is 35 will retire. They will be
replaced by five new members whose mean age is 24. What will the new
mean age of the Comets be ?
Answer #1

Solution :

Given that,

The mean age of the 30 members of the Ft. Wayne Comets is 28

so, 30 * 28 = 840

After the current season, five members whose mean age is 35 will retire then

5 * 35 = 175

840 - 175 = 665

They will be replaced by five new members whose mean age is 24,

5 * 24 = 120

So, the new mean age of the Comets be ,

New age = (665 + 120) / 30 = 785 / 30 = 26.167